OpenStudy (ivancsc1996):

Why, in the Euler Lagrange equation isn't d/dt of the partial derivative of L with respect to q' actually the partial derivativ of L with respect to q' and t? Remember that L is L(t,q,q')

OpenStudy (ivancsc1996):

Otherwise stated. Why is the equation:\[\frac{ d }{ dt} L _{q'}-L _{t}=0\]instead of:\[L _{tq'}-L _{t}=0\]

OpenStudy (kmeis002):

Small type, but the Lagrangian is: \[ \frac{\partial L}{\partial q} - \frac{d}{dt}\frac{\partial L}{\partial q'} \] L is a function of two functions of t, q and q', but we only care about the explicit functions and not the implied t when deriving. The time derivative arises from derivative when integrating to minimize \[ ( \frac{\partial L}{\partial q} \delta q+ \frac{\partial L}{\partial q'} \delta q')\] but note that q is a only a function of t so \[ q'(t) = \frac{dq}{dt}\] Using this, we obtain \[ ( \frac{\partial L}{\partial q} \delta q+ \frac{\partial L}{\partial q'} \frac{d}{dt}\ \delta q)\] And more math is done to simplify to our lagrangian and set equal to 0.